# On higher-dimensional symmetric designs

### v1

@@ 4. Difference sets
This is a special case of [5, Theorem 2.9]. Properties of these difference cubes were studied in Section 3 of [21]. In Section 4, the construction was generalized to the so-called group cubes [21, Theorem 4.1]. This construction gives examples not equivalent with any difference cube, although it still requires difference sets. The only known examples of $\mathcal{C}$-cubes not coming from [21, Theorem 4.1] are the $\mathcal{C}^{3}(16,6,2)$-cubes of [21, Proposition 5.3].
@@
is an OA-representation $\overline{C}$ of a cube $C\in\mathcal{P}^{n}(v,k,\lambda)$ indexed by $G$. Properties of these cubes were studied in Section 3 of [22]. There it was shown that $n$-dimensional difference sets can be normalized so that all $n$-tuples in $D$ start with a 00 coordinate. We will now prove a linear bound on the dimension of $D$ using a different kind of normalization.
@@
For a group element $g\in G$ and an $n$-tuple $d=(d_{1},\ldots,d_{n})\in G^{n}$, denote by $g+_{x}d=(d_{1},\ldots,d_{x-1},g+d_{x},d_{x+1},\ldots,d_{n})$. We claim that $D^{\prime}=\{g+_{x}d\mid d\in D\}$ is also an $n$-dimensional $(v,k,\lambda)$ difference set. For any other index $y\in\{1,\ldots,n\}$, the set $\{d_{x}-d_{y}\mid d\in D^{\prime}\}=\{g+d_{x}-d_{y}\mid d\in D\}$ is a left translate of the difference set $\{d_{x}-d_{y}\mid d\in D\}$, hence also a $(v,k,\lambda)$ difference set. By repeated application of this transformation we can make an $n$-dimensional difference set $D^{\prime}$ containing the $n$-tuple $(0,\ldots,0)$. Now any other $n$-tuple $d\in D^{\prime}$ must have distinct coordinates, because $d_{x}=d_{y}$ would imply that the set of differences $\{d_{x}-d_{y}\mid d\in D^{\prime}\}$ contains fewer than $k$ elements of $G$. Therefore, the number of coordinates $n$ cannot exceed $v=|G|$. ∎
@@
The largest integer $n$ such that an $n$-dimensional $(v,k,\lambda)$ difference set in $G$ exists was denoted by $\mu_{G}(v,k,\lambda)$ in [22]. To determine values of this function, a computer classification of small $n$-dimensional difference sets was performed using the library of difference sets available in the GAP package [29]. In [22], calculations were performed in GAP. Exact values of $\mu_{G}(16,6,2)$ were determined in [22, Table 3] for $10$ of the $14$ groups of order $16$, and lower bounds were given for the remaining $4$ groups.
@@
The previously known values of $\mu$ are given in [22, Tables 2 and 3] and can also be read from Table 4. The groups of order $16$ with IDs $1$, $7$, $12$, and $14$ are missing from Table 4. The former two groups are the cyclic group $\mathbb{Z}_{16}$ and the dihedral group $D_{16}$, which do not contain difference sets. For the latter two groups we could not completely classify $n$-dimensional difference sets due to large numbers of inequivalent $\mathcal{P}$-cubes arising from them. Example 4.3 establishes an improved lower bound $\mu_{G}(16,6,2)\geq 11$ for the group with ID $12$. The group with ID $14$ is the elementary abelian group $\mathbb{Z}_{2}^{4}$ and will be dealt with in Corollary 4.5.
Example 4.3.
[…]
We see that equality is reached quite often in the upper bound $\mu_{G}(v,k,\lambda)\leq v$ of Theorem 4.1. For parameters $(q,(q-1)/2,(q-3)/4)$ this is explained by the construction of $q$-dimensional Paley difference sets in the additive group of $\mathbb{F}_{q}$, $q\equiv 3\pmod{4}$; see [22, Theorem 4.1]. It was noted that the same construction works for cyclotomic difference sets ($4$th and $8$th powers in $\mathbb{F}_{q}$ for appropriate orders $q$), but this does not explain the equality $\mu(13,4,1)=\mu(13,9,6)=13$. We will now generalize this construction to difference sets with arbitrary parameters in elementary abelian groups.
@@
Led $D=\{d_{1},\ldots,d_{k}\}\subseteq G$ be a $(q,k,\lambda)$ difference set, $\alpha\in\mathbb{F}_{q}$ a primitive element, and $w=(0,1,\alpha,\alpha^{2},\ldots,\alpha^{q-2})\in G^{q}$. An extension of $D$ to dimension $q$ is $\vec{D}=\{d_{1}w,\ldots,d_{k}w\}$, where each coordinate of $w$ is multiplied by an element of $D$. To prove that $\vec{D}$ is indeed a $q$-dimensional $(q,k,\lambda)$ difference set, we have to check that the differences of any two coordinates are difference sets in $G$. The first coordinate of any vector in $\vec{D}$ is 00, and if we subtract the coordinate with $\alpha^{y}$ in $w$, we get $\{-\alpha^{y}d_{1},\ldots,-\alpha^{y}d_{k}\}$. This is a $(q,k,\lambda)$ difference set in $G$ because it is the image of $D$ by the group automorphism $\varphi:G\to G$, $\varphi(g)=-\alpha^{y}g$. Next we subtract two non-zero coordinates, say with $\alpha^{x}$ and $\alpha^{y}$ in $w$ for some $x<y$. We get the set
@@
If $G$ is abelian, automorphisms and antiautomorphisms coincide. Antiautomorphisms of non-abelian groups are the functions $-\varphi$, where $\varphi$ is an automorphism. The crucial fact in the proof of Theorem 4.4 is that the image $\varphi(D)$ of a $(v,k,\lambda)$ difference set $D$ by an automorphism $\varphi$ is also a $(v,k,\lambda)$ difference set. The same holds true for antiautomorphisms of nonabelian groups.
@@
Theorem 4.4 can be seen as a special case of Theorem 4.7 by using multiplication in $\mathbb{F}_{q}$. The functions $\varphi_{x}:\mathbb{F}_{q}\to\mathbb{F}_{q}$, $\varphi_{x}(g)=\alpha^{x}g$ for $x=0,\ldots,q-2$ constitute a regular set of automorphism of the additive group, yielding the construction of Theorem 4.4. Kerdock sets provide more examples of regular sets in elementary abelian $2$-groups [18, 3, 7, 15, 16] and $3$-groups [28].
Theorem 4.7 also applies to groups that are not elementary abelian. For example, $\varphi_{1}(g)=g$ and $\varphi_{2}(g)=2g$ are automorphisms of the group $\mathbb{Z}_{15}$. The difference $(\varphi_{1}-\varphi_{2})(g)=-g$ is an automorphism as well. Thus, $\{\varphi_{1},\varphi_{2}\}$ is a regular set and all $(15,7,3)$ and $(15,8,4)$ difference sets extend at least to dimension $3$. In Table 4 we see that there is a $4$-dimensional $(15,8,4)$ difference set, but it cannot be obtained from Theorem 4.7 because $\mathbb{Z}_{15}$ does not allow regular sets of size $3$. The next result applies to difference sets in arbitrary cyclic groups $\mathbb{Z}_{v}$.
@@
Let $p$ be the smallest prime divisor of $v$. Then any cyclic $(v,k,\lambda)$ difference set extends to dimension $p$.
@@
For a non-abelian example we turn to groups of order $27$. There are five such groups, but only two of them contain nontrivial difference sets [29]. These are the elementary abelian group $\mathbb{Z}_{3}\times\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ (GAP group ID 5) and a semidirect product $\mathbb{Z}_{9}\rtimes\mathbb{Z}_{3}$ (GAP group ID 4). Difference sets in the former group extend to dimension $27$ by Theorem 4.4. A regular set in the latter group is given in the next example, using multiplicative notation.
@@
By Theorem 4.7, all $(27,13,6)$ and $(27,14,7)$ difference sets in $G$ extend to dimension $3$.
@@ 5. Final observations
A curious fact visible in Table 4 is equality of the numbers for some complementary parameters. Ordinary ($2$-dimensional) difference sets and designs are rarely discussed for both sets of complementary parameters, because complementation is a bijection. This is also true for higher-dimensional $\mathcal{C}$-cubes, but not for $\mathcal{P}$-cubes as we have seen in Proposition 2.3. However, the numbers in Table 4 coincide for complementary Paley-Hadamard parameters $\mathcal{P}^{n}(4m-1,2m-1,m-1)$ and $\mathcal{P}^{n}(4m-1,2m,m)$ when $4m-1=p^{s}$ is a prime power and $G=(\mathbb{Z}_{p})^{s}$ is an elementary abelian group. For small parameters, this can be explained by the following bijection.
Every $(4m-1,2m-1,m-1)$ difference set $D\subseteq G$ can be uniquely extended to a $(4m-1,2m,m)$ difference set $D\cup\{a\}$ by adding the element $a=-2\sum_{d\in D}d$. We have checked that this is a bijection between $(4m-1,2m-1,m-1)$ and $(4m-1,2m,m)$ difference sets in the groups $\mathbb{Z}_{7}$, $\mathbb{Z}_{11}$, $\mathbb{Z}_{19}$, and $\mathbb{Z}_{23}$ from Table 4 and also in $G=(\mathbb{Z}_{3})^{3}$. We don’t have a proof for arbitrary elementary abelian groups. However, if this is indeed a bijection between ordinary difference sets, then it is easy to establish a $1$-to-$1$ correspondence between $n$-dimensional difference sets. Given an $n$-dimensional $(4m-1,2m-1,m-1)$ difference set $D\subseteq G^{n}$, one simply adds the $n$-tuple $\left(-2\sum_{d\in D}d_{1},\ldots,-2\sum_{d\in D}d_{n}\right)$ to obtain an $n$-dimensional $(4m-1,2m,m)$ difference set.
If a $(4m-1,2m-1,m-1)$ difference set $D\subseteq G$ extends to a $(4m-1,2m,m)$ difference set $D\cup\{a\}$, then the complement $D^{\prime}=G\setminus(D\cup\{a\})$ is also a $(4m-1,2m-1,m-1)$ difference set in $G$. Then $(D-a)\cup(D^{\prime}-a)\cup\{0\}$ is a tiling of $G$ by two difference sets, and $D-a$ is a skew Hadamard difference set by [4, Theorem 8]. A long-standing conjecture was that the classical Paley difference sets are the only skew Hadamard difference sets in elementary abelian groups, but it has been refuted in [8, 9] for groups of orders $3^{m}$, $m\geq 5$ odd. It might also be the case that our bijection holds up only for elementary abelian groups of small orders, but we don’t have a counterexample.
Yet another apparent symmetry in Table 4 is equality of the numbers for “complementary dimensions” $n$ and $v-n$. Equality holds for some small parameters $(v,k,\lambda)$ when the upper bound of Theorem 4.1 is reached: $(7,3,1)$, $(7,4,2)$, $(11,5,2)$, $(11,6,3)$, and $(13,4,1)$. It breaks for the complementary parameters $(13,9,6)$, while for $(19,9,4)$, $(19,10,5)$, $(23,11,5)$, and $(23,12,6)$ equality holds except for dimension $n=3$. For $(31,6,1)$ it breaks for $n=3$, $4$, $5$, $6$ and holds for the remaining dimensions. We don’t have an explanation for this phenomenon.

### v2

@@ 4. Difference sets
This is a special case of [5, Theorem 2.9]. Properties of these difference cubes were studied in Section 3 of [21]. In Section 4, the construction was generalized to the so-called group cubes [21, Theorem 4.1]. This construction gives examples not equivalent to any difference cube, having non-isomorphic $(v,k,\lambda)$ designs as $2$-sections. The construction of [21, Theorem 4.1] still requires difference sets; the only known examples of $\mathcal{C}$-cubes not coming from this constructions are the $\mathcal{C}^{3}(16,6,2)$-cubes of [21, Proposition 5.3].
@@
is an OA-representation $\overline{C}$ of a cube $C\in\mathcal{P}^{n}(v,k,\lambda)$ indexed by $G$. Properties of these cubes were studied in Section 3 of [22]. There, it was shown that $n$-dimensional difference sets can be normalized so that all $n$-tuples in $D$ start with a 00 coordinate. Using a different kind of normalization, we will now prove a bound on the dimension of $D$ stronger than the bound (2).
@@
For a group element $g\in G$ and an $n$-tuple $d=(d_{1},\ldots,d_{n})\in G^{n}$, denote by $g+_{x}d=(d_{1},\ldots,d_{x-1},g+d_{x},d_{x+1},\ldots,d_{n})$. We claim that $D^{\prime}=\{g+_{x}d\mid d\in D\}$ is also an $n$-dimensional $(v,k,\lambda)$ difference set. For any other index $y\in\{1,\ldots,n\}$, the set $\{d_{x}-d_{y}\mid d\in D^{\prime}\}=\{g+d_{x}-d_{y}\mid d\in D\}$ is a left translate of the difference set $\{d_{x}-d_{y}\mid d\in D\}$, hence also a $(v,k,\lambda)$ difference set. By repeated application of this transformation, we can make an $n$-dimensional difference set $D^{\prime}$ containing the $n$-tuple $(0,\ldots,0)$. Now any other $n$-tuple $d\in D^{\prime}$ must have distinct coordinates, because $d_{x}=d_{y}$ would imply that the set of differences $\{d_{x}-d_{y}\mid d\in D^{\prime}\}$ contains fewer than $k$ elements of $G$. Therefore, the number of coordinates $n$ cannot exceed $v=|G|$. ∎
@@
The largest integer $n$ such that an $n$-dimensional $(v,k,\lambda)$ difference set in $G$ exists was denoted by $\mu_{G}(v,k,\lambda)$ in [22]. To determine values of this function, a computer classification of small $n$-dimensional difference sets was performed using the library of difference sets available in the GAP package DifSets [29]. In [22], calculations were performed in GAP. Exact values of $\mu_{G}(16,6,2)$ were determined in [22, Table 3] for $10$ of the $14$ groups of order $16$, and lower bounds were given for the remaining $4$ groups.
@@
The previously known values of $\mu$ are given in [22, Tables 2 and 3] and can also be read from Table 4. The groups of order $16$ with IDs $1$, $7$, $12$, and $14$ are missing from Table 4. The former two groups are the cyclic group $\mathbb{Z}_{16}$ and the dihedral group $D_{16}$, which do not contain difference sets. For the latter two groups we could not completely classify $n$-dimensional difference sets due to large numbers of inequivalent $\mathcal{P}$-cubes arising from them. Example 4.3 establishes an improved lower bound $\mu_{G}(16,6,2)\geq 11$ for the group with ID $12$. The group with ID $14$ is the elementary abelian group $\mathbb{Z}_{2}^{4}$, and will be dealt with in Corollary 4.5.
Example 4.3.
[…]
We see that equality is reached quite often in the upper bound $\mu_{G}(v,k,\lambda)\leq v$ of Theorem 4.1. For parameters $(q,(q-1)/2,(q-3)/4)$, this is explained by the construction of $q$-dimensional Paley difference sets in the additive group of $\mathbb{F}_{q}$, $q\equiv 3\pmod{4}$; see [22, Theorem 4.1]. It was noted that the same construction works for cyclotomic difference sets ($4$th and $8$th powers in $\mathbb{F}_{q}$ for appropriate orders $q$), but this does not explain the equality $\mu(13,4,1)=\mu(13,9,6)=13$. We will now generalize this construction to difference sets with arbitrary parameters in elementary abelian groups. The key observation in the proof of the following theorem is that group automorphisms $\varphi:G\to G$ preserve difference sets in $G$.
@@
Let $D=\{d_{1},\ldots,d_{k}\}\subseteq G$ be a $(q,k,\lambda)$ difference set, $\alpha\in\mathbb{F}_{q}$ a primitive element, and $w=(0,1,\alpha,\alpha^{2},\ldots,\alpha^{q-2})\in G^{q}$. An extension of $D$ to dimension $q$ is $\vec{D}=\{d_{1}w,\ldots,d_{k}w\}$, where each coordinate of $w$ is multiplied by an element of $D$. To prove that $\vec{D}$ is indeed a $q$-dimensional $(q,k,\lambda)$ difference set, we have to check that the differences of any two coordinates are difference sets in $G$. The first coordinate of any vector in $\vec{D}$ is 00, and if we subtract the coordinate with $\alpha^{y}$ in $w$, we get $\{-\alpha^{y}d_{1},\ldots,-\alpha^{y}d_{k}\}$. This is a $(q,k,\lambda)$ difference set in $G$ because it is the image of $D$ by the group automorphism $\varphi:G\to G$, $\varphi(g)=-\alpha^{y}g$. Next, we subtract two non-zero coordinates, say with $\alpha^{x}$ and $\alpha^{y}$ in $w$ for some $x<y$. We get the set
@@
If $G$ is abelian, automorphisms and antiautomorphisms coincide. Antiautomorphisms of non-abelian groups are the functions $-\varphi$, where $\varphi$ is an automorphism. Thus, antiautomorphisms of $G$ also map $(v,k,\lambda)$ difference set $D\subseteq G$ to difference sets with the same parameters.
@@
Theorem 4.4 can be seen as a special case of Theorem 4.7 by using multiplication in $\mathbb{F}_{q}$. The functions $\varphi_{x}:\mathbb{F}_{q}\to\mathbb{F}_{q}$, $\varphi_{x}(g)=\alpha^{x}g$ for $x=0,\ldots,q-2$ constitute a regular set of automorphisms of the additive group, yielding the construction of Theorem 4.4. Kerdock sets provide more examples of regular sets in elementary abelian $2$-groups [3, 7, 15, 16, 18] and $3$-groups [28].
Theorem 4.7 also applies to groups that are not elementary abelian. For example, $\varphi_{1}(g)=g$ and $\varphi_{2}(g)=2g$ are automorphisms of the group $\mathbb{Z}_{15}$. The difference $(\varphi_{1}-\varphi_{2})(g)=-g$ is an automorphism as well. Thus, $\{\varphi_{1},\varphi_{2}\}$ is a regular set and all $(15,7,3)$ and $(15,8,4)$ difference sets extend at least to dimension $3$. In Table 4, we see that there is a $4$-dimensional $(15,8,4)$ difference set, but it cannot be obtained from Theorem 4.7 because $\mathbb{Z}_{15}$ does not allow regular sets of size $3$. The next result applies to difference sets in arbitrary cyclic groups $\mathbb{Z}_{v}$.
@@
Let $p$ be the smallest prime divisor of $v$. Then any cyclic $(v,k,\lambda)$ difference set extends at least to dimension $p$.
@@
For a non-abelian example, we turn to groups of order $27$. There are five such groups, but only two of them contain nontrivial difference sets [29]. These are the elementary abelian group $\mathbb{Z}_{3}\times\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ (GAP group ID 5) and a semidirect product $\mathbb{Z}_{9}\rtimes\mathbb{Z}_{3}$ (GAP group ID 4). Difference sets in the former group extend to dimension $27$ by Theorem 4.4. A regular set in the latter group is given in the next example, using multiplicative notation.
@@
By Theorem 4.7, all $(27,13,6)$ and $(27,14,7)$ difference sets in $G=\langle a,b\mid a^{9}=b^{3}=1,\,ba=a^{4}b\rangle$ extend to dimension $3$.
@@ 5. Final observations
A curious fact visible in Table 4 is equality of the numbers for some complementary parameters. Ordinary ($2$-dimensional) difference sets and designs are rarely discussed for both sets of complementary parameters, because complementation is a bijection. This is also true for higher-dimensional $\mathcal{C}$-cubes, but not for $\mathcal{P}$-cubes as we have seen in Proposition 2.3. However, the numbers in Table 4 coincide for complementary Paley-Hadamard parameters $\mathcal{P}^{n}(4m-1,2m-1,m-1)$ and $\mathcal{P}^{n}(4m-1,2m,m)$, when $4m-1=p^{s}$ is a prime power and $G=(\mathbb{Z}_{p})^{s}$ is an elementary abelian group. For small parameters, this can be explained by the following bijection.
Every $(4m-1,2m-1,m-1)$ difference set $D\subseteq G$ can be uniquely extended to a $(4m-1,2m,m)$ difference set $D\cup\{a\}$ by adding the element $a=-2\sum_{d\in D}d$. We have checked that this is a bijection between $(4m-1,2m-1,m-1)$ and $(4m-1,2m,m)$ difference sets in the groups $\mathbb{Z}_{7}$, $\mathbb{Z}_{11}$, $\mathbb{Z}_{19}$, and $\mathbb{Z}_{23}$ from Table 4, and also in $G=(\mathbb{Z}_{3})^{3}$. We don’t have a proof for arbitrary elementary abelian groups. However, if this is indeed a bijection between ordinary difference sets, then it is easy to establish a $1$-to-$1$ correspondence between $n$-dimensional difference sets. Given an $n$-dimensional $(4m-1,2m-1,m-1)$ difference set $D\subseteq G^{n}$, one simply adds the $n$-tuple $\left(-2\sum_{d\in D}d_{1},\ldots,-2\sum_{d\in D}d_{n}\right)$ to obtain an $n$-dimensional $(4m-1,2m,m)$ difference set.
If a $(4m-1,2m-1,m-1)$ difference set $D\subseteq G$ extends to a $(4m-1,2m,m)$ difference set $D\cup\{a\}$, then the complement $D^{\prime}=G\setminus(D\cup\{a\})$ is also a $(4m-1,2m-1,m-1)$ difference set in $G$. Then $(D-a)\cup(D^{\prime}-a)\cup\{0\}$ is a tiling of $G$ by two difference sets, and $D-a$ is a skew Hadamard difference set by [4, Theorem 8]. A long-standing conjecture was that the classical Paley difference sets are the only skew Hadamard difference sets in elementary abelian groups, but it has been refuted in [8, 9] for groups of orders $3^{m}$, $m\geq 5$ odd. It might also be the case that our bijection holds only for elementary abelian groups of small orders, but we don’t have a counterexample.
Yet another apparent symmetry in Table 4 is equality of the numbers for “complementary dimensions” $n$ and $v-n$. Equality holds for some small parameters $(v,k,\lambda)$ when the upper bound of Theorem 4.1 is reached: $(7,3,1)$, $(7,4,2)$, $(11,5,2)$, $(11,6,3)$, and $(13,4,1)$. It breaks for the complementary parameters $(13,9,6)$, while for $(19,9,4)$, $(19,10,5)$, $(23,11,5)$, and $(23,12,6)$ equality holds except for dimension $n=3$. For $(31,6,1)$, it breaks for $n=3$, $4$, $5$, $6$ and holds for the remaining dimensions. We don’t have an explanation for this phenomenon.